Oxidation number of iodine in IO3,IO4,KI and I2 respectively are____________.

A

1,1,0,+1

B

+3,+5,+7,0

C

+5,+7,1,0

D

1,5,1,0

E

2,5,1,0

 

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
Oxidation number of oxygen is−2andKis+1.InIO3−​,oxidation number of iodine is(x+3(−2)=−1,x=+5)+5.
InIO4−​,oxidation number of iodine is(x+4(−2)=−1,x=+7)+7.
InKI,oxidation number of iodine is(1+x=0,x=−1)−1.
InI2​,oxidation number of iodine is0.